Circular No. 6728

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970828 D. Smith, A. Levine, R. Remillard, and A. Wood, Massachusetts Insti-

tute of Technology, communicate: “A refined analysis of the RXTE ASM observations of GRB 970828 (described on IAUC 6726), using the two 90-s observations simultaneously, yields a best-fit location and 90-percent confidence ellipse. The center of this ellipse lies at the position reported on IAUC 6726; its axes lie along the local α and δ lines, such that its major axis spans 5′.0 in α and its minor axis 2′.0 in δ. This 90-percent-confidence ellipse includes an estimate of the systematic errors added in quadrature, and it is smaller than the 1σ error region reported previously.”

K. Hurley, Space Sciences Laboratory, Berkeley, on behalf of the Ulysses Gamma-Ray Burst Team; C. Kouveliotou, G. Fishman, C. Meegan, and V. Connaughton, on behalf of the Gamma Ray Observatory BATSE Team; and J. van Paradijs, University of Amsterdam and University of Alabama in Huntsville, report: “A preliminary triangulation has been done using Ulysses and BATSE data. The resulting position is an annulus whose halfwidth is conservatively estimated to be 0^{o}.027 at 3σ confidence, centered at α = 10^{h}46^{m}13^{s}, δ = +19^{o}10′.3 (equinox 2000.0), with central radius 83^{o}.518. This annulus passes through the corners of the RXTE error box (IAUC 6726) but does not reduce its size substantially. The annulus is therefore also consistent with the RXTE PCA source reported on IAUC 6727. Further processing will reduce the annulus width.”

SUPERNOVA 1997de IN NGC 6769 G. J. Garradd, Loomberah, N.S.W., reports the following precise posi-

tion for SN 1997de (cf. IAUC 6725), obtained with a 0.25-m f/4.1 reflector (+ CCD) on Aug. 29.444 UT: α = 19^{h}18^{m}20^{s}.82, δ = −60^{o}29′42′′.0 (equinox 2000.0); this corresponds to an offset 14′′.2 west and 20′′.7 north of the center of NGC 6769. The estimated magnitude of the supernova was 17.5.

COMET C/1997 O1 (TILBROOK) Visual m_{1} estimates: Aug. 4.4 UT, 11.0: (F. Farrell, Christies Beach,

S. Australia, 0.15-m reflector); 7.38, 10.3 (J. Seach, Mclean, N.S.W., 0.25-m reflector); 7.88, 11.1: (J. Carvajal, Avila, Spain, 0.44-m reflector); 21.41, 11.0 (D. A. J. Seargent, Cowra, N.S.W., 25×100 binoculars).

1997 August 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green